OpenStudy (anonymous):

There are 1,000 fish in a lake. Each year the population declines by 15%, but afterward, the lake is restocked with 500 additional fish. Write a recursive rule to find the number of fish in the lake for any year.

OpenStudy (anonymous):

0.15*1000=150+500=650

OpenStudy (anonymous):

hope i am right

OpenStudy (anonymous):

i think it wants a recursion

OpenStudy (anonymous):

are telling me that i am wrong?

OpenStudy (anonymous):

you telling*

OpenStudy (anonymous):

\[A_1=1000, A_n=.85\times A_{n-1}+500\]might work to decrease a number by \(15\%\) you multiply it by \(100\%-15\%=85\%=.85\) then add \(500\)